OpenStudy (anonymous):

solve each problem. Make sure your answer is not a improper fraction. a.) 6/7x7=. b.) 2x3/5=. c.) 9x4/7=. d.) 2/3x6=.

OpenStudy (texaschic101):

6/7 x 7 (if it makes it easier for you, write it like this: 6/7 x 7/1 (now just multiply across) 42/7 -- reduce 7/1, or just 7 ==================== 2/1 x 3/5 = 6/5 -- change to mixed number 1 1/5 (one and one- fifth) ==================== 9/1 x 4/7 = 36/7 -- change to mixed number 5 6/7 (five and six-sevenths ==================== 2/3 x 6/1 = 12/3 -- reduce 4/1, or just 4

OpenStudy (anonymous):

maybe im wrong but my answer for c.) is 5 1/7 check out that please

OpenStudy (texaschic101):

oops...you are correct blanca. Good catch. C is 5 1/7

OpenStudy (anonymous):

I think for a.) is 6 check out this too please

OpenStudy (texaschic101):

it is....a. is 6, I am sorry....I have the flu and have just taken some nyquil and it is kicking my butt. My brain is kind of in la-la land. I probably just need to go to bed.

OpenStudy (anonymous):

no problem and thank you for your help I hope you feel better